OpenStudy (anonymous):

What is the simplified form of the quantity 15 x y squared over x squared minus 5x plus 6 all over the quantity 5 x squared y over 2x squared minus 7x plus 3 ?

OpenStudy (campbell_st):

so the question is \[\frac{\frac{15xy^2}{x^2 - 5x + 6}}{\frac{5x^2y}{2x^2 - 7x + 3}}\] so when you divide by a fraction, flip the 2nd fraction and multiply \[\frac{5xy^2}{x^2 - 5x + 6} \times \frac{2x^2 - 7x + 3}{5x^2y}\] I would factor the quadractics eliminate any common factors what is left should be the answer

OpenStudy (anonymous):

wow i didnt even think about keep change flip!